For the function f(x)=(x+4)/(x^2-16)
a) Find the factored form
b) Find any vertical asymptotes
c) Find any holes
d) Find the intercepts
e) Find the horizontal asymptote

Answer by lwsshak3(11628)

(a) f(x) =(x+4)/(x^2-16) =(x+4)/(x+4)(x-4)=1/(x-4)
(b) vertical asymptotes, x=4
(c) hole in function at x=-4
(d) x-intercept, none; y-intercept,-1/4
(e) horizontal asymptote, y=0 or x-axis
note: function is undefined at x=4 and not defined at x=-4